# Recursive subhomogeneity of orbit-breaking subalgebras of $\mathrm{C}^{*}$-algebras associated to minimal homeomorphisms twisted by line bundles

### v1

@@ 4. Orbit-breaking algebras
Let $A$ be a $\mathrm{C}^{*}$-algebra and $\mathcal{E}$ a $\mathrm{C}^{*}$-correspondence over $A$ with the structure map $\varphi_{\mathcal{E}}$. We say that $\mathcal{E}$ is minimal if, for any non-zero ideal $J\subset A$, if $\langle\mathcal{E},\varphi_{E}(J)\mathcal{E}\rangle_{\mathcal{E}}\subset J$, then $J=A$. This definition is a generalization of the corresponding definition for a single homeomorphism $\alpha:X\to X$ on a compact metric space: We say that $\alpha$ is minimal if, for every non-empty closed subset $Y\subset X$, if $\alpha(Y)\subset Y$, then $Y=X$.
@@
    If $\alpha$ ia minimal, then $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is simple if and only if $Y\cap\alpha^{n}(Y)=\emptyset$ for every $n\in\mathbb{Z}\setminus\{0\}$.
@@
That $\Gamma(\mathscr{V},\alpha)$ is minimal if and only if $\alpha$ is minimal if and only if the crossed product $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$ is simple follows from [2, Proposition 3.9, Corollary 3.11]. Suppose that $\alpha$ is minimal. That $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is simple if and only if $Y\cap\alpha^{n}(Y)\neq\emptyset$ for $n\neq 0$ follows from [24]. (Note that this is a slight improvement of the result stated in [2]). Applying [23, Theorem 7.3, Proposition 8.8] to Cuntz–Pimsner algebras $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$ and $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ yields (3). ∎
@@ 5. Recursive subhomogeneous algebras
Let $X$ be an infinite compact metric space, $\alpha:X\to X$ a minimal homeomorphism, $\mathscr{V}$ a line bundle over $X$ and $Y\subset X$ a closed subset with non-empty interior. As the construction of a recursive subhomogeneous decomposition of an orbit-breaking algebra $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is quite technical, this section recalls the definition of a recursive subhomogeneous algebras as well as the recursive subhomogeneous algebra structure of orbit-breaking algebras for crossed products by minimal homeomorphisms. We then describe the conceptual steps in obtaining the recursive subhomogeneous of $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$, which highlights the similarities and differences to the construction in the crossed product case.
@@
A recursive subhomogeneous decomposition for a given $\mathrm{C}^{*}$-algebra is in general not unique. Observe that any recursive subhomogeneous $\mathrm{C}^{*}$-algebra can be written in the form
$R=\left[\cdots\left[\left[C_{0}\oplus_{C^{(0)}_{1}}C_{1}\right]\oplus_{C_{2}^{(0)}}\right]\cdots\right]\oplus_{C_{l}^{(0)}}C_{l},$ (5)
where $C_{j}=C(X_{j},M_{n(j)})$ for compact Hausdorff spaces $X_{j}$, $n(j)\in\mathbb{Z}_{>0}$ and $C^{(0)}_{j}=C(X_{j}^{(0)},M_{n(j)})$ for (possibly empty) compact subsets $X^{(0)}_{j}\subset X_{j}$, and where the maps $C_{j}\to C_{j}^{(0)}$ are the restriction maps. The expression (5) is called a recursive subhomogeneous decomposition of $R$.
@@
    $R^{(k)}=\left[\cdots\left[\left[C_{0}\oplus_{C^{(0)}_{1}}C_{1}\right]\oplus_{C_{2}^{(0)}}\right]\cdots\right]\oplus_{C_{j}^{(0)}}C_{j},$
@@
    $X_{0},\dots,X_{l}$ are called the base spaces of the decomposition and their disjoint union $\bigsqcup_{j=1}^{l}X_{j}$ is called the total space,
@@
Let $\alpha$ be a minimal homeomorphism on a compact metrizable space $X$ and $Y$ be a closed subset of $X$ with non-empty interior as before. In the following, we will denote the orbit-breaking algebra $\mathrm{C}^{*}(C(X),C_{0}(X\setminus Y)u)$ by $A_{Y}$. For $x\in Y$, the first return time $r_{Y}(x)$ of $x$ to $Y$ is the smallest integer $n\geq 1$ such that $\alpha^{n}(x)\in Y$. Then it can be checked that $\sup_{x\in Y}r_{Y}(x)<\infty$. Let $r_{1}<\dots<r_{K}$ be the distinct values of the first return times to Y and set
$Y_{k}=\{x\in Y\mid r_{Y}(x)=r_{k}\},$ (6)
for all $1\leq k\leq K$. If $Y$ is understood, we use $r(x)$ in place of $r_{Y}(x)$. Note that $\cup_{i=1}^{k}Y_{i}$ is always closed for every $k=1,\dots,K$, and
$Y=\bigsqcup_{k=1}^{K}Y_{k},\quad\mbox{and}\quad X=\bigsqcup_{k=1}^{K}\bigsqcup_{i=0}^{r_{k}-1}\alpha^{i}(Y_{k}).$ (7)
@@
$Y_{k},\alpha(Y_{k}),\cdots,\alpha^{r_{k}-1}(Y_{k})$
@@
If $X$ is a Cantor set, then each $Y_{k}$ is clopen and compact. Set $p_{k}=\chi_{Y_{k}}$, then each $p_{k}$ commutes with elements in $A_{Y}$ and we have $A_{Y}=\bigoplus_{k=1}^{K}p_{k}A_{Y}p_{k}$. Then $\pi_{k}\colon p_{k}A_{Y}p_{k}\to C(Y_{k},M_{r_{k}})$ given by
$\pi_{k}(f)=\mathrm{diag}(f|_{Y_{k}},\dots,f\circ\alpha^{r_{k}-1}|_{Y_{k}})$
@@
$\pi_{k}(\chi_{\alpha^{j}(Y_{k})}u\chi_{\alpha^{j-1}(Y_{k})})=e_{j,j-1}\otimes 1.$
Then each $\pi_{k}$ is an isomorphism, specifically, $A_{Y}\cong\bigoplus_{k=1}^{K}C(Y_{k},M_{r_{k}})$.
If $X$ is not a Cantor set, then $Y_{k}$’s are not necessarily closed and boundaries $\alpha^{j}(\overline{Y_{k}})$ can overlap which lead to a recursive subhomogeneous decomposition for orbit breaking subalgebras instead of direct limit of homogeneous algebras.
Set
@@
Let $\pi_{k}:A_{Y}\to C(\overline{Y_{k}},M_{r_{k}})$ be the composition of $\pi$ with the projection into the $k$-summand of $B_{Y}$. Set $\pi^{(k)}=\oplus_{i=1}^{k}\pi_{k}$ and $A_{k}=\pi^{(k)}(A_{Y})$. We say that an element $b=(b_{1},\dots,b_{K})$ has a boundary decomposition property if it satisfies the following: for any $x\in\overline{Y_{k}}\setminus Y_{k}$, $k=1,\dots,K$ with
@@
$b_{k}(x)=\mathrm{diag}(b_{t_{1}}(x),b_{t_{2}}(\alpha^{r_{t_{1}}}(x)),\dots,b_{t_{s}}(\alpha^{r_{t_{1}}+\dots+r_{t{s-1}}}(x))).$
Then $b=(b_{1},\dots,b_{K})\in B_{Y}$ is in $\pi(A_{Y})$ if and only if $b$ has the boundary decomposition property. This result enables us to explicitly describe the recursive subhomogeneous construction of orbit-breaking subalgebra $A_{Y}$. Fix $1\leq k\leq K$ to define $\varphi_{k}:A_{k-1}\to C(\overline{Y_{k}}\setminus Y_{k},M_{r_{k}})$, take arbitrary $(b_{1},\dots,b_{k-1})\in A_{k-1}$. By the definition there is $b=(b_{1},\dots,b_{k-1},b_{k},\dots,b_{K})\in\pi(A_{Y})$. Since $b$ satisfies the boundary decomposition property, $b_{k}|_{\overline{Y_{k}}\setminus Y_{k}}$ is only depend on $(b_{1},\dots,b_{k-1})$. Now, define $\varphi_{k}(b_{1},\dots,b_{k-1})(x)=b_{k}(x)$ and let $\rho_{k}:C(\overline{Y_{k}},M_{r_{k}})\to C(\overline{Y_{k}}\setminus Y_{k},M_{r_{k}})$ denote the restriction map. Then
$A_{k}=A_{k-1}\oplus_{C(\overline{Y_{k}}\setminus Y_{k},M_{r_{k}})}C(\overline{Y_{k}},M_{r_{k}}).$
@@
$[\dots[C(Y_{1},M_{r_{1}})\oplus_{C(\overline{Y_{2}}\setminus Y_{2},M_{r_{2}})}C(\overline{Y}_{2},M_{r_{2}})]\dots]\oplus_{C(\overline{Y_{k}}\setminus Y_{k},M_{r_{k}})}C(\overline{Y_{k}},M_{r_{k}})]$ (8)
Now let $\mathscr{V}$ be a line bundle over $X$ and set $\mathcal{E}:=\Gamma(\mathscr{V},\alpha)$. We describe the conceptual steps in obtaining a recursive subhomogeneous decomposition for the orbit-breaking algebra $\mathcal{O}(\mathcal{E}_{Y})$.
@@
Step 1. In Section 6, we begin with introducing the building blocks in the recursive subhomogeneous decomposition the of orbit-breaking algebra $\mathcal{O}(\mathcal{E}_{Y})$ Set $\mathscr{D}^{(r_{k})}=\mathscr{V}^{(0)}\oplus\dots\oplus\mathscr{V}^{(r_{k}-1)}$ where $\mathscr{V}^{(i)}$ is the line bundle defined in (3), and let $\mathscr{M}_{k}:=End(\mathscr{D}^{(r_{k})})|_{\overline{Y}_{k}}$ be the endomorphism bundle (see Section 6). We show in Theorem 8.7 that $C(\overline{Y}_{k},M_{r_{k}})$ in (8) will be replaced by $\Gamma(\mathscr{M}_{k})$ when the twist is not trivial.
@@
Step 2. In Section 7, we construct an embedding from the orbit-breaking algebra $\mathcal{O}(\mathcal{E}_{Y})$ into $\Gamma(\mathscr{M}_{1})\oplus\dots\oplus\Gamma(\mathscr{M}_{K})$. For this, we introduce a covariant representation $(\pi_{k},\tau_{k})$ of the Hilbert $C(X)$-bimodule $\mathcal{E}_{Y}$ to $\Gamma(\mathscr{M}_{k})$ which induces a *-homomorphism, again denoted by $\pi_{k}$, from the orbit-breaking algebra to $\Gamma(\mathscr{M}_{k})$, for all $1\leq k\leq K$. In Proposition 7.4 we show that $\pi=\oplus_{k=1}^{K}\pi_{k}$ is injective and its image admits a gauge action, which, by the gauge invariant uniqueness theorem, tells us that $\pi$ is a -isomorphism from $\mathcal{O}(\mathcal{E}_{Y})$ onto its image.
@@
Step 3. To describe the pullbacks in the RSH structure, we need to understand the image of the *-representation $\pi$. We generalize the boundary decomposition for elements in $\Gamma(\mathscr{M}_{1})\oplus\dots\oplus\Gamma(\mathscr{M}_{K})$ in Definition 8.1. Then by theorem 8.6 the range of $\pi$ is precisely the set of all elements with boundary decomposition property.
@@
Step 4. Finally we obtain the main result of the paper. The boundary decomposition property of elements in the image of $\pi$ enables us to define *-homomorphisms $\varphi_{k}:(\oplus_{i=1}^{k-1}\pi_{i})(\mathcal{O}(\Gamma(\mathcal{E}_{Y}))\to\Gamma(\mathscr{M}|_{\overline{Y}_{k}\setminus Y_{k}})$. Then by Theorem 8.7, the pullback $(\oplus_{i=1}^{k-1}\pi_{i})(\mathcal{O}(\mathcal{E}_{Y}))\oplus_{\Gamma(\mathscr{M}_{k}|_{\overline{Y}_{k}\setminus Y_{k}})}\Gamma(\mathscr{M}_{k})$ via $\varphi_{k}$ and the restriction map is equal to $(\oplus_{i=1}^{k}\pi_{i})(\mathcal{O}(\mathcal{E}_{Y}))$. That is,
Every $\Gamma(\mathscr{M}_{k})$ is itself an RSH algebra [32, Proposition 1.7], which gives us the RSH decomposition of the orbit-breaking subalgebra.

### v2

@@ 4. Orbit-breaking algebras
Let $A$ be a $\mathrm{C}^{*}$-algebra and $\mathcal{E}$ a $\mathrm{C}^{*}$-correspondence over $A$ with the structure map $\varphi_{\mathcal{E}}$. We say that $\mathcal{E}$ is minimal if, for any non-zero ideal $J\subset A$, if $\langle\mathcal{E},\varphi_{\mathcal{E}}(J)\mathcal{E}\rangle_{\mathcal{E}}\subset J$, then $J=A$. This definition is a generalization of the corresponding definition for a single homeomorphism $\alpha:X\to X$ on a compact metric space: We say that $\alpha$ is minimal if, for every non-empty closed subset $Y\subset X$, if $\alpha(Y)\subset Y$, then $Y=X$.
@@
    If $\alpha$ is minimal, then $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is simple if and only if $Y\cap\alpha^{n}(Y)=\emptyset$ for every $n\in\mathbb{Z}\setminus\{0\}$.
@@
That $\Gamma(\mathscr{V},\alpha)$ is minimal if and only if $\alpha$ is minimal if and only if the crossed product $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$ is simple follows from [2, Proposition 3.9, Corollary 3.11]. Suppose that $\alpha$ is minimal. That $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is simple if and only if $Y\cap\alpha^{n}(Y)\neq\emptyset$ for $n\neq 0$ follows from [26]. (Note that this is a slight improvement of the result stated in [2]). Applying [25, Theorem 7.3, Proposition 8.8] to Cuntz–Pimsner algebras $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$ and $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ yields (3). ∎
@@ 5. Recursive subhomogeneous algebras
In this section, we begin by recalling the definition of and terminology for recursive subhomogeneous algebras.
Next, we recall the recursive subhomogeneous algebra structure of orbit-breaking subalgebras for crossed products by minimal homeomorphisms. This is a special case of our construction.
We then provide the conceptual steps for our more general orbit-breaking algebras, $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ for $X$ an infinite compact metric space, $\alpha:X\to X$ a minimal homeomorphism, $\mathscr{V}$ a line bundle over $X$ and $Y\subset X$ a closed subset with non-empty interior. The aim is to highlight the similarities and differences to the construction in the crossed product case, as well as to aid the reader in understanding the technical constructions in subsequent sections.
5.1. Recursive subhomogeneous algebras
@@
Observe that any recursive subhomogeneous $\mathrm{C}^{*}$-algebra $R$ can be written in the form
$R\cong\left[\cdots\left[\left[C_{0}\oplus_{C^{(0)}_{1}}C_{1}\right]\oplus_{C_{2}^{(0)}}\right]\cdots\right]\oplus_{C_{l}^{(0)}}C_{l},$ (5)
where $C_{j}=C(X_{j},M_{n(j)})$ for compact Hausdorff spaces $X_{j}$, $n(j)\in\mathbb{Z}_{>0}$ and $C^{(0)}_{j}=C(X_{j}^{(0)},M_{n(j)})$ for (possibly empty) compact subsets $X^{(0)}_{j}\subset X_{j}$, and where the maps $C_{j}\to C_{j}^{(0)}$ are the restriction maps. The expression (5) is called a recursive subhomogeneous decomposition of $R$. A recursive subhomogeneous decomposition for a given $\mathrm{C}^{*}$-algebra is in general not unique.
@@
    $R^{(k)}=\left[\cdots\left[\left[C_{0}\oplus_{C^{(0)}_{1}}C_{1}\right]\oplus_{C_{2}^{(0)}}\right]\cdots\right]\oplus_{C_{k}^{(0)}}C_{k},$
@@
    $X_{0},\dots,X_{l}$ are called the base spaces of the decomposition and their disjoint union $\bigsqcup_{j=0}^{l}X_{j}$ is called the total space,
@@
5.2. Orbit-breaking in crossed products by homeomorphisms
Let $\alpha$ be a minimal homeomorphism on an infinite compact metrizable space $X$ and let $Y$ be a closed subset of $X$ with non-empty interior. For $x\in Y$, the first return time $r_{Y}(x)$ of $x$ to $Y$ is the smallest integer $n\geq 1$ such that $\alpha^{n}(x)\in Y$. Since $\alpha$ is minimal and $Y$ has non-empty interior, $\sup_{x\in Y}r_{Y}(x)<\infty$. A compactness argument shows that there are only finitely many distinct return times to $Y$, which we denote by $r_{1}<\dots<r_{K}$. For $1\leq k\leq K$, set
$Y_{k}:=\{x\in Y\mid r_{Y}(x)=r_{k}\}.$ (6)
For every $k=1,\dots,K$, the union $\cup_{i=1}^{k}Y_{i}$ is closed. It is easy to check that
$Y=\bigsqcup_{k=1}^{K}Y_{k},\quad\text{and}\quad X=\bigsqcup_{k=1}^{K}\bigsqcup_{i=0}^{r_{k}-1}\alpha^{i}(Y_{k}).$ (7)
@@ 5.2. Orbit-breaking in crossed products by homeomorphisms
$Y_{k},\alpha(Y_{k}),\,\dots,\,\alpha^{r_{k}-1}(Y_{k})$
@@
Denote by $A_{Y}$ the orbit-breaking subalgebra $\mathrm{C}^{*}(C(X),C_{0}(X\setminus Y)u)\subset C(X)\rtimes_{\alpha}\mathbb{Z}$. Here $u$ denotes the unitary implementing $\alpha$ in the crossed product.
5.2.1. Orbit-breaking for minimal Cantor systems
Suppose that $X$ is the Cantor space. In [37], Putnam constructs Rokhlin towers for $Y$ consisting of clopen sets. Since the sets are clopen, their indicator functions are continuous and generate projections in $A_{Y}$. Since $\alpha^{j}(Y_{k})\cap Y=\emptyset$ for every $11\leq j\leq r_{k}-$ 1 we have that $\chi_{\alpha^{j}(Y_{k})}u\in A_{Y}$ for every $11\leq j\leq r_{k}-$ 1. Setting $e_{j,j-1}:=\chi_{\alpha^{j}(Y_{k})}u$, it is easy to check that $e_{j,j-1}$ generate a copy of $M_{r_{k}}$. This allows Putnam to show that $A_{Y}$ is isomorphic with a direct sum of homogeneous algebras $M_{r_{k}}(Y_{k})$ [37, Theorem 3.3].
Example 5.3 ([37, Theorem 3.3], see also [18, Lemma 11.2.22]).
Let $X$ be a Cantor set, $Y$ be a non-empty compact open subset and $\alpha:X\to X$ be a minimal homeomorphism. Then each $Y_{k}$ is clopen and compact. Set $p_{k}:=\chi_{Y_{k}}$. Then each $p_{k}$ commutes with elements in $A_{Y}$ and we have $A_{Y}=\bigoplus_{k=1}^{K}p_{k}A_{Y}p_{k}$. Then $\pi_{k}\colon p_{k}A_{Y}p_{k}\to C(Y_{k},M_{r_{k}})$ given by
$\pi_{k}(f)=\mathrm{diag}(f|_{Y_{k}},\dots,f\circ\alpha^{r_{k}-1}|_{Y_{k}}),$
@@ 5.2.1. Orbit-breaking for minimal Cantor systems
$\pi_{k}(\chi_{\alpha^{j}(Y_{k})}u)=e_{j,j-1}\otimes 1$
is an isomorphism and we have $A_{Y}\cong\bigoplus_{k=1}^{K}C(Y_{k},M_{r_{k}})$.
5.2.2. Orbit-breaking subalgebras of arbitrary crossed products
If $X$ is not a Cantor space and $Y$ is not clopen, then the sets $Y_{k}$’s are not necessarily closed. If $x\in\overline{Y_{k}}\setminus Y_{k}$, then by continuity, $\alpha^{r_{k}}(y)\in Y$. However, this might not be the first return time of $y$ to $Y$. In particular, for any $1<k\leq K$, the intersection $\overline{Y_{k}}\cup Y_{j}$ for $j<k$ may be nonempty. In this case, we can only expect a recursive subhomogeneous decomposition for $A_{Y}$, rather than a direct sum of homogeneous algebras.
Example 5.4 ([30], see also [18, Theorem 11.3.19]).
Let $X$ be a compact metric space, $\alpha:X\to X$ a minimal homeomorphism, and $Y$ a closed subset of $X$ with non-empty interior. Set
@@ 5.2.2. Orbit-breaking subalgebras of arbitrary crossed products
Let $\pi_{k}:A_{Y}\to C(\overline{Y_{k}},M_{r_{k}})$ be the composition of $\pi$ with the projection onto the $k$th summand of $B_{Y}$. Set $\pi^{(k)}=\oplus_{i=1}^{k}\pi_{k}$ and $B_{k}=\pi^{(k)}(A_{Y})$.
We say that an element $b=(b_{1},\dots,b_{K})$ has the boundary decomposition property if it satisfies the following: for any $x\in\overline{Y_{k}}\setminus Y_{k}$, $k=1,\dots,K$, with
@@
$b_{k}(x)=\mathrm{diag}(b_{t_{1}}(x),b_{t_{2}}(\alpha^{r_{t_{1}}}(x)),\dots,b_{t_{s}}(\alpha^{r_{t_{1}}+\dots+r_{t{m-1}}}(x))).$
An element $b=(b_{1},\dots,b_{K})\in B_{Y}$ is in $\pi(A_{Y})$ if and only if $b$ has the boundary decomposition property. This result enables us to explicitly describe the recursive subhomogeneous construction of the orbit-breaking algebra $A_{Y}$.
The $1$st stage algebra is given by $B_{1}=C(\overline{Y_{1}},M_{r_{1}})$. (In fact, $Y_{1}$ is already a closed subset.)
Fix $1\leq k\leq K$ and let $B_{k-1}$ denote the $(k-1)$th stage algebra. To define the $k$th stage algebra $B_{k}$, we construct the pullback
[…]
where $\rho_{k}$ is restriction, and $\varphi_{k}:B_{k-1}\to C(\overline{Y_{k}}\setminus Y_{k},M_{r_{k}})$ is defined as follows: Let $(b_{1},\dots,b_{k-1})\in B_{k-1}$. For $k\leq j\leq K$, there are $b_{j}\in C(\overline{Y_{j}},M_{r_{j}})$ such that $b=(b_{1},\dots,b_{k-1},b_{k},\dots,b_{K})\in\pi(A_{Y})$. Since $b$ satisfies the boundary decomposition property, $b_{k}|_{\overline{Y_{k}}\setminus Y_{k}}$ only depends on $(b_{1},\dots,b_{k-1})$. Now, define $\varphi_{k}(b_{1},\dots,b_{k-1})(x)=b_{k}(x)$. Then
$B_{k}=B_{k-1}\oplus_{C(\overline{Y_{k}}\setminus Y_{k},M_{r_{k}})}C(\overline{Y_{k}},M_{r_{k}}).$
@@
$[\cdots[C(Y_{1},M_{r_{1}})\oplus_{C(\overline{Y_{2}}\setminus Y_{2},M_{r_{2}})}C(\overline{Y}_{2},M_{r_{2}})]\dots\oplus_{C(\overline{Y_{K}}\setminus Y_{K},M_{r_{K}})}C(\overline{Y_{K}},M_{r_{K}})].$ (8)
5.3. Generalizing to a non-trivial line bundle
Now assume further that there is a line bundle $\mathscr{V}$ over $X$ and set $\mathcal{E}:=\Gamma(\mathscr{V},\alpha)$. The main goal of this paper is to show that, for $Y\subset X$ a closed subset with non-empty interior, the orbit-breaking algebra $\mathcal{O}(\mathcal{E}_{Y})$ is a recursive subhomogeneous algebra. Here we outline the steps involved in the construction.
@@ 5.3. Generalizing to a non-trivial line bundle
Step 1. In Section 6, we introduce the building blocks in the recursive subhomogeneous decomposition of the orbit-breaking algebra $\mathcal{O}(\mathcal{E}_{Y})$. Let $\mathscr{D}^{(r_{k})}:=\mathscr{V}^{(0)}\oplus\dots\oplus\mathscr{V}^{(r_{k}-1)}$ denote the rank $r_{k}$ vector bundle over $X$ given by the direct sum of the line bundles $\mathscr{V}^{(i)}$ defined in (3), and let $\mathscr{M}_{k}:=End(\mathscr{D}^{(r_{k})})|_{\overline{Y}_{k}}$ be the endomorphism bundle (see Section 6). When the twist $\mathscr{V}$ is no longer trivial, we replace $C(\overline{Y}_{k},M_{r_{k}})$ in (8) by $\Gamma(\mathscr{M}_{k})$, see Theorem 8.7.
@@
Step 2. In Section 7, we construct an embedding from the orbit-breaking algebra $\mathcal{O}(\mathcal{E}_{Y})$ into $\Gamma(\mathscr{M}_{1})\oplus\dots\oplus\Gamma(\mathscr{M}_{K})$. For this, we introduce a covariant representation $(\pi_{k},\tau_{k})$ of the Hilbert $C(X)$-bimodule $\mathcal{E}_{Y}$ to $\Gamma(\mathscr{M}_{k})$ which induces a *-homomorphism, again denoted by $\pi_{k}$, from the orbit-breaking algebra to $\Gamma(\mathscr{M}_{k})$, for all $1\leq k\leq K$. In Proposition 7.5 we show that $\pi=\oplus_{k=1}^{K}\pi_{k}$ is injective and its image admits a gauge action, which, by the gauge invariant uniqueness theorem, implies that $\pi$ is a -isomorphism from $\mathcal{O}(\mathcal{E}_{Y})$ onto its image.
@@
Step 3. To describe the pullbacks in the RSH structure, we need to understand the image of the *-representation $\pi$. We generalize the boundary decomposition property for elements in $\Gamma(\mathscr{M}_{1})\oplus\dots\oplus\Gamma(\mathscr{M}_{K})$ in Definition 8.1. Then by Theorem 8.6 the range of $\pi$ is precisely the set of all elements with boundary decomposition property.
@@
Step 4. Finally we obtain the main result of the paper. The boundary decomposition property of elements in the image of $\pi$ enables us to define *-homomorphisms
$\varphi_{k}:(\oplus_{i=1}^{k}\pi_{i})(\mathcal{O}(\mathcal{E}_{Y}))\to\Gamma(\mathscr{M}_{k+1}|_{\overline{Y_{k+1}}\setminus Y_{k+1}}),$
$k=1,\dots,K-1$. Then by Theorem 8.7, the pullback
$(\oplus_{i=1}^{k}\pi_{i})(\mathcal{O}(\mathcal{E}_{Y}))\oplus_{\Gamma(\mathscr{M}_{k+1}|_{\overline{Y_{k+1}}\setminus Y_{k+1}})}\Gamma(\mathscr{M}_{k})$
via $\varphi_{k}$ and the restriction map $\rho_{k}:\Gamma(\mathscr{M}_{k+1})\to\Gamma(\mathscr{M}_{k+1}|_{\overline{Y_{k+1}}\setminus Y_{k+1}})$ is equal to $(\oplus_{i=1}^{k+1}\pi_{i})(\mathcal{O}(\mathcal{E}_{Y}))$. That is,
$\mathcal{O}(\mathcal{E}_{Y})\cong$ $\ \pi(\mathcal{O}(\mathcal{E}_{Y}))$
$=$ $\ [\cdots[\Gamma(\mathscr{M}_{1})\oplus_{\Gamma(\mathscr{M}_{2}|_{\overline{Y_{2}}\setminus Y_{2}})}\Gamma(\mathscr{M}_{2})]\oplus\cdots]\oplus_{\Gamma(\mathscr{M}_{K}|_{\overline{Y_{K}}\setminus Y_{K}})}\Gamma(\mathscr{M}_{K}).$
Since the section algebra of each matrix bundles has a recursive subhomogeneous decomposition [35, Proposition 1.7], we conclude that the iterated pullback describing the orbit-breaking algebra also has an RSH decomposition.